Broussard Skateboard's sales are expected to increase by 20% from $8.2 million in 2016 to $9.84 million in 2017. Its assets totaled $6 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 3%, and the forecasted payout ratio is 75%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar.
$

Assume that an otherwise identical firm had $7 million in total assets at the end of 2016. The identical firm's capital intensity ratio (A0*/S0) is -Select-higher than lower than equal to

Item 2 than Broussard's; therefore, the identical firm is -Select- less more the same

Item 3 capital intensive - it would require -Select-a smaller a larger the same

Item 4 increase in total assets to support the increase in sales.
Answer #1
1] AFN is given by the equation:
AFN = (A/S) x (Δ Sales) - (L/S) x (Δ Sales) - (PM x FS x (1-d))
where,
A / S = Assets that change directly with sales.
Δ sales = Change in sales
L / S = Liabilities that change directly with sales
PM = Profit Margin on Sales = net income / sales.
FS = Forecasted Sales
d = dividend payout ratio
(1 - d) = retention ratio
Substituting required values in the above equation we have,
AFN = (6000000/8200000) x 1640000 - (900000/8200000) x 1640000 - 9840000*3%*(1-75%) [ $             946,200
The identical firm's capital intensity ratio (A0*/S0) is higher than Broussard's.
It would require a larger increase in total assets
to support the increase in sales.